OpenStudy (anonymous):

help please!!! A company's monthly profit increases by $1,000 each month. In January, the profit of the company was $25,000. If x = 0 represents January, which of the following equations represents the profit as a function of time (in months)? A. y = 25,000x + 1,000 B. y = 1,000x C. y = 1,000x – 25,000 D. y = 1,000x + 25,000

OpenStudy (♪chibiterasu):

I would definitely say D, x being 0 just makes y = 25,000, which is the monthly profit.

OpenStudy (anonymous):

Profit can increase from an increase in revenue, a decrease in expenses or a combination of the two.

OpenStudy (anonymous):

@♪Chibiterasu

OpenStudy (♪chibiterasu):

What does this have to do with the question?

OpenStudy (anonymous):

can you help me in this one ?

OpenStudy (anonymous):

Rewrite the standard form equation below into slope-intercept form. x – 2y = 3

OpenStudy (♪chibiterasu):

Ok, well first we have to get the x onto the other side. Do you know how? Slope-intercept is: y = mx + b

OpenStudy (anonymous):

so its like y=3/2x-1/2

OpenStudy (anonymous):

or y= (-)3/2x-1/2

OpenStudy (♪chibiterasu):

Not exactly . . . First subtract x from both sides: x – x -2y = -x + 3 -2y = -x + 3

OpenStudy (♪chibiterasu):

Now divide -2 from both sides: -2y/-2 = -x/-2 + 3 y = -1/2x + 3

OpenStudy (anonymous):

i got it thanks

OpenStudy (♪chibiterasu):

-x/-2 is the same as -1/2x

OpenStudy (♪chibiterasu):

Ok :3

OpenStudy (anonymous):

can you help me in another one?

OpenStudy (♪chibiterasu):

Sure. o:

OpenStudy (anonymous):

Which of the following equations is in standard form? A. y=1/2x B. y= 1/2x-3 C. 4y-1=1/2x-3 D. 2x3y=14

OpenStudy (anonymous):

is b

OpenStudy (♪chibiterasu):

Standard form = Ax + By = C

OpenStudy (♪chibiterasu):

@desmonds No, and don't blurt out the answer please.

OpenStudy (anonymous):

ok

OpenStudy (♪chibiterasu):

So from the example, can you tell which one is correct? Now we know it isn't B. :P

OpenStudy (anonymous):

b?

OpenStudy (anonymous):

can you help me too

OpenStudy (♪chibiterasu):

No o:

OpenStudy (♪chibiterasu):

A and B are replaced with digits. ;o

OpenStudy (anonymous):

or D

OpenStudy (♪chibiterasu):

Yep!! :3

OpenStudy (anonymous):

can you help me

OpenStudy (anonymous):

Which of the following equations is an example of the general form of the equation of a line in standard form? A. y = mx + b B. Ax = By + C C. By = Ax + C D. Ax + By = C

OpenStudy (♪chibiterasu):

Post your question in the "Ask a question..." box

OpenStudy (♪chibiterasu):

@desmonds

OpenStudy (anonymous):

so on this one is B

OpenStudy (anonymous):

Which of the following equations is an example of the general form of the equation of a line in standard form? A. y = mx + b B. Ax = By + C C. By = Ax + C D. Ax + By = C

OpenStudy (anonymous):

standard form

OpenStudy (♪chibiterasu):

Not exactly, the equal symbol is placed incorrectly.

OpenStudy (♪chibiterasu):

@desmonds Please delete that and post it in the Ask a question... box, it's rude to post your question in someone else's. :(

OpenStudy (anonymous):

its A then

OpenStudy (♪chibiterasu):

No, that's slope-intercept. Ax + By = C ;o

OpenStudy (anonymous):

so its A y = mx + b

OpenStudy (♪chibiterasu):

No . . .

OpenStudy (anonymous):

B

OpenStudy (♪chibiterasu):

It's D :|

OpenStudy (♪chibiterasu):

Ax + By = C

OpenStudy (anonymous):

ok thanks

OpenStudy (♪chibiterasu):

Np :P